Fulbright Corp. uses the periodic inventory system. During its first year of operations, Fulbright made the following purchases (listed in chronological order of acquisition):

  • 40 units at $109 per unit
  • 73 units at $80 per unit
  • 174 units at $62 per unit


Sales for the year totaled 271 units, leaving 16 units on hand at the end of the year.


Ending inventory using the LIFO method is:

Multiple Choice

  • $1,170.

  • $1,744.

  • $1,042.

  • $992.

Fulbright Corp. uses the periodic inventory system. During its first year of operations, Fulbright made the following purchases (listed in chronological order of acquisition):

  • 59 units at $157 per unit
  • 165 units at $172 per unit
  • 189 units at $152 per unit

Sales for the year totaled 365 units, leaving 48 units on hand at the end of the year.

In comparing the ending inventory balances of FIFO and LIFO, the ending inventory value under FIFO less the ending inventory balance under LIFO results in a difference of:

Multiple Choice

  • $53.

  • $(240).

  • $0.

  • $240.

Inventory records for Herb's Chemicals revealed the following:


March 1, 2021, inventory: 1,500 gallons @ $8.20 per gallon = $12,300

Purchases: Sales:
Mar. 10 700 gals @ $ 8.25 Mar. 5 500 gals
Mar. 16 900 gals @ $ 8.30 Mar. 14 800 gals
Mar. 23 700 gals @ $ 8.35 Mar. 20 600 gals
Mar. 26 800 gals


Ending inventory assuming LIFO in a periodic inventory system would be:

Multiple Choice

  • $9,020.

  • $9,135.

  • $9,075.

  • $9,035.

Alison's dress shop buys dresses from McGuire Manufacturing. Alison purchased dresses from McGuire on July 17 and received an invoice with a list price amount of $5,800 and payment terms of 2/10, n/30. Alison uses the net method to record purchases. Alison should record the purchase at:

Multiple Choice

  • $2,842.

  • $5,800.

  • $5,916.

  • $5,684.

The use of LIFO during a long inflationary period can result in:

Multiple Choice

  • Significant cash flow advantages over FIFO.

  • An inflated balance sheet.

  • A net increase in income tax expense.

  • A reduction in inventory turnover over FIFO.
Answer #1
1 Fulbright Corp.
Answer: $1,744
Under the LIFO method, the last purchased units will be issued first leaving the earliest purchased in the ending inventory.
Ending inventory = 16 units x $109 = $1744
2 Fulbright Corp.
Answer: $(240)
Ending inventory (FIFO) = 48 x $152 = $7296
Ending inventory (LIFO) = 48 x $157 = $7536
Ending inventory (FIFO) - Ending inventory (LIFO) = $7296 - $7536 = $(240)
3 Herb's Chemicals
Answer: $9,020
Under the LIFO method, the last purchased units will be issued first leaving the earliest purchased in the ending inventory.
Ending inventory (units) = Beginning inventory + Purchases - Sales = 1500 + (700 + 900 + 700) - (500 + 800 + 600 + 800) = 1500 + 2300 - 2700 = 1100
Ending inventory (LIFO) = 1100 x $8.20 = $9020
4 Alison's Dress Shop
Answer: $5,684
Under the net method of recording purchases, the purchases will be recorded net of the purchase discount.
Discount = 2% x $5800 = $116
Net purchases recorded = $5800 - $116 = $5684

Per HOMEWORKLIB RULES the first 4 MCQs have been answered. Please post any remaining separately. Thank you.